# Scaling Laws for Black-box Adversarial Attacks

### v3

@@ Abstract
Adversarial examples usually exhibit good cross-model transferability, enabling attacks on black-box models with limited information about their architectures and parameters, which are highly threatening in commercial black-box scenarios. Model ensembling is an effective strategy to improve the transferability of adversarial examples by attacking multiple surrogate models. However, since prior studies usually adopt few models in the ensemble, there remains an open question of whether scaling the number of models can further improve black-box attacks. Inspired by the scaling law of large foundation models, we investigate the scaling laws of black-box adversarial attacks in this work. Through theoretical analysis and empirical evaluations, we conclude with clear scaling laws that using more surrogate models enhances adversarial transferability. Comprehensive experiments verify the claims on standard image classifiers, diverse defended models and multimodal large language models using various adversarial attack methods. Specifically, by scaling laws, we achieve 90%+ transfer attack success rate on even proprietary models like GPT-4o. Further visualization indicates that scaling laws exists on the interpretability and semantics of adversarial perturbations.
@@ 1 Introduction
Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [54]. However, deep neural networks (DNNs) remain susceptible to adversarial examples [35, 73] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions. Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [21], assessing model robustness [98], and informing the development of more secure algorithms [65].
@@
Figure 1: Key observation of scaling laws for the transfer-based black-box attack. By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a clear scaling law.
@@
Adversarial attacks are typically categorized into white-box and black-box attacks based on whether the attacker has knowledge of the architecture and parameters of the target model. Given the greater risks with commercial services, e.g., GPT-4 [68], black-box attack methodologies [60, 22] have attracted growing attention. One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models. Researchers have devised advanced optimization algorithms [65, 81, 57] and data augmentation techniques [63, 23] in transfer-based attacks. Model ensembling [60, 22] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples. By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [22, 9].
@@
Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [9, 97, 94]. Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [48]. Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [22, 9], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide an asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models including adversarial training [65, 86] and DiffPure [67] (Sec. 3.3), and modern multimodal large language models (Sec. 3.4 and Sec. 4) using various attack methods (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer-based attacks.
More concretely, through scaling laws, we achieve over 30% improvement against vision classifiers and over 10% improvement against defended models. By extending the surrogate models to a group of CLIP [70] vision encoders, we achieve over 90% attack success rate against GPT-4o and GPT-4-Turbo, and over 80% against other commercial MLLMs. This highlights the vulnerabilities of these models to perturbations and the significance of scaling laws for ensemble attacks. We also demonstrate clear scaling laws on the interpretability and semantics of adversarial perturbations. Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than designing intricate algorithms for a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy MLLMs.
@@ 2 Backgrounds
In this section, we provide necessary backgrounds, including problem formulation and algorithms.
2.1 Problem Formulation
In this work, we focus on the challenging targeted attack setting. Let $\mathcal{F}$ denote the set of all image classifiers. Given a natural image $\bm{x}_{\text{nat}}$ and target class $\bm{y}$, the goal of targeted transfer-based attacks is to craft an adversarial example $\bm{x}$ that is misclassified to $\bm{y}$ by all models in $\mathcal{F}$. It can be formulated as a constrained optimization problem:
$\min_{\bm{x}}\mathbb{E}_{f\in\mathcal{F}}\left[\mathcal{L}(f(\bm{x}),\bm{y})\right],\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (1)
where $\mathcal{L}$ is the loss function (e.g., cross-entropy loss), and we consider adversarial perturbations in $\ell_{\infty}$ norm.
Generally, we take a limited set of “training” classifiers $\mathcal{F}_{T}:=\{f_{i}\}_{i=1}^{T}\subset\mathcal{F}$ as model ensemble to approximate problem (1) and therefore the optimization objective becomes:
$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (2)
where $T=\left\lvert\mathcal{F}_{T}\right\rvert$ is the cardinality of model ensemble.
@@ 2.2 Prerequisite Algorithms
Researchers have been developing sophisticated algorithms for ensemble attacks, and we select the state-of-the-art algorithms from them to conduct our experiments.
MI-FGSM [22] is a pioneering method of transfer-based attacks. It stabilizes the gradient update directions by integrating a momentum term into the iteration:
@@
and updates $\bm{x}$ by $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$.
Common Weakness Method (CWA) [9] seeks to identify common vulnerabilities across surrogate models by encouraging flatter loss landscapes and closer proximity between the local optima of these models. It minimizes the second-order Taylor series:
@@
where $\bm{p}_{i}$ denotes the closest optimum of model $f_{i}$ to $\bm{x}$ and $H_{i}$ is the Hessian matrix at $\bm{p}_{i}$.
Spectrum Simulation Attack (SSA) [63] applies a spectrum transformation $\mathcal{T}$ to the $\bm{x}$ and calculate gradients with respective to $\mathcal{T}(\bm{x})$ instead of $\bm{x}$, and can be combined with CWA to form SSA-CWA algorithm (pseudocode is provided in Appendix C). As shown in Table 1, SSA-CWA gives the best attack success rate. Therefore, in this paper, we use SSA-CWA as the core algorithm. We also explore the scaling laws of other algorithms in Sec. 5.1.
ASR Target: Car Target: Ship
Methods MI-CWA SSA-CWA MI-CWA SSA-CWA
Swin-v2 0.44 0.65 0.36 0.62
ConvNeXt 0.52 0.58 0.53 0.62
ResNeXt50 0.69 0.79 0.66 0.77
Table 1: Comparison of attack success rate (ASR) between SSA and MI algorithms. We follow the experiment settings in Sec. 3.2, and measure the ASR on 2 target classes and 3 black-box models.
Figure 2: Comparisons between CWA and MI-FGSM methods. We draw inputs $\bm{x}$ randomly from the dataset in Sec. 3.2 and measure the averaged gradient magnitudes on $\bm{x}$ after 10 steps of iterations. The results are recorded across $T$ from $1$ to $30$. On the right side, there are the visualizations of gradients of $T=30$.
Mitigating Gradient Conflict. We discover a new phenomenon of gradient conflict when ensembling a large number of surrogates. As shown in Fig. 2, as the number of models increases, the conflict between gradients of different models using common optimizers (e.g., MI-FGSM) becomes more severe, eventually reducing the ensemble gradient to zero and resulting in complete optimization stagnation. We conjecture this is because the gradients of normally trained models commonly lack semantics and have random directions; thus, naïve ensemble methods that directly average these gradients lead to strong gradient conflicts, significantly impacting optimization. To address this problem, we suggest to use optimizers that reduce the variance of gradients (e.g., SVRE [92]) or encourage the cosine similarity of gradients (e.g., CWA [9]). As demonstrated in Fig. 2, these methods fully avoid this issue. We analyze the gradient conflict on different attackers in more detail in Sec. 5.1.
@@ 3 Scaling Laws for Adversarial Attacks
In this section, we present our key findings on scaling laws for black-box adversarial attacks. Sec. 3.1 introduces fundamental concepts and theorems that give us motivations for the study of scaling laws. Sec. 3.2 contains our experiment settings and derivations of empirical scaling laws. Sec. 3.3 demonstrates clear scaling laws when attacking common defended models. Sec. 3.4 shows the results over multimodal large language models.
3.1 Motivations
@@
[…]
@@
[…]
@@
[…]
Theorem 2 gives us motivations for further exploring the effects of ensemble cardinality $T$ as empirical loss converges with a factor of $T$. This leads us to recovering the scaling laws on black-box adversarial attacks.
3.2 Scaling Laws on Image Classifiers

### v4

@@ Abstract
Adversarial examples exhibit cross-model transferability, enabling threatening black-box attacks on commercial models. Model ensembling, which attacks multiple surrogate models, is a known strategy to improve this transferability. However, prior studies typically use small, fixed ensembles, which leaves open an intriguing question of whether scaling the number of surrogate models can further improve black-box attacks. In this work, we conduct the first large-scale empirical study of this question. We show that by resolving gradient conflict with advanced optimizers, we discover a robust and universal log-linear scaling law through both theoretical analysis and empirical evaluations: the Attack Success Rate (ASR) scales linearly with the logarithm of the ensemble size $T$. We rigorously verify this law across standard classifiers, SOTA defenses, and MLLMs, and find that scaling distills robust, semantic features of the target class. Consequently, we apply this fundamental insight to benchmark SOTA MLLMs. This reveals both the attack’s devastating power and a clear robustness hierarchy: we achieve 80%+ transfer attack success rate on proprietary models like GPT-4o, while also highlighting the exceptional resilience of Claude-3.5-Sonnet. Our findings urge a shift in focus for robustness evaluation: from designing intricate algorithms on small ensembles to understanding the principled and powerful threat of scaling.
@@ 1 Introduction
Over the past decade, deep learning has achieved remarkable advancements across various computer vision tasks, leading to its widespread applications [lecun2015deep]. However, deep neural networks (DNNs) remain susceptible to adversarial examples [goodfellow2014explaining, szegedy2013intriguing] – perturbations that are subtle and nearly imperceptible but can significantly mislead the model, resulting in incorrect predictions. Adversarial attacks have garnered considerable interest due to their implications for understanding DNN functionality [dong2017towards], assessing model robustness [zhang2024benchmarking], and informing the development of more secure algorithms [madry2017towards].
@@
Figure 1: Key observation of scaling laws for the transfer-based black-box attack. By scaling the number of surrogate models in the ensemble, we observe that the attack success rate and the discriminative loss on the target model follow a log-linear scaling law.
@@
Adversarial attacks are typically categorized into white-box and black-box attacks based on whether the attacker has knowledge of the architecture and parameters of the target model. Given the greater risks with commercial services, e.g., GPT-4 [openai2024gpt4technicalreport], black-box attack methodologies [liu2016delving, dong2018boosting] have attracted growing attention. One prominent approach within black-box attacks is the transfer-based attack, which approximately optimizes the adversarial examples against the black-box model by minimizing the discriminative loss on a selection of predefined surrogate models. Researchers have devised advanced optimization algorithms [madry2017towards, wang2021enhancing, lin2019nesterov] and data augmentation techniques [long2022frequency, dong2019evading] in transfer-based attacks. Model ensembling [liu2016delving, dong2018boosting] further facilitates these methods by leveraging multiple surrogate models simultaneously to improve the transferability of adversarial examples. By averaging the gradients or predictions from different models, ensemble attack methods seek to increase the transferability across various target models [dong2018boosting, chen2023rethinking].
@@
Despite the progress in ensemble attack algorithms, the number of surrogate models used in previous research is often limited [chen2023rethinking, zhang2024does, yao2025understandingmodelensembletransferable]. Recent advancement in large language models shows that scaling training data can significantly improve the generalization ability [kaplan2020scalinglawsneurallanguage]. Since the transfer-based attack is also a generalization problem that expects adversarial examples from surrogate models to generalize to target models, while the training data are surrogate models rather than text or images [dong2018boosting, chen2023rethinking], this equivalence prompts an intriguing question: Can we achieve significant improvements in transfer-based attacks by scaling the number of surrogate models, akin to the scaling of data and model parameters in large language models like GPT?
In this work, we conduct the first large-scale empirical investigation into the scaling laws of black-box adversarial attacks, given that the gradient conflict issue is properly resolved (Sec. 2.2). Our central finding is the discovery of a robust and universal log-linear scaling law: the Attack Success Rate (ASR) on black-box targets scales linearly with the logarithm of the ensemble cardinality $T$. We provide an asymptotic analysis on the convergence rate of generation error and adversarial examples under idealized IID assumptions (Sec. 3.1). Our core contribution, however, is the empirical discovery and verification of the precise log-linear form of these scaling laws (Sec. 3.3). We rigorously demonstrate the universality of these scaling laws across diverse target families (Sec. 3.4), and provide a qualitative explanation for this phenomenon (Sec. 5.2): as $T$ increases, the optimizer is forced to discard model-specific, non-robust features and instead learns the robust, semantic features of the target class, resulting in highly transferable perturbations.
Inspired by this fundamental insight, we then extend our scaling laws to attack the foundational vision encoders (e..g., CLIP) of modern MLLMs (Sec. 4) to demonstrate its practical power as both a SOTA attack and a novel robustness benchmark. We achieve devastatingly high ASR on top-tier models, including 90.0% ASR on Qwen3-VL-235B and 85.0% ASR on GPT-4o. Meanwhile, our scaling benchmark reveals a stark hierarchy in SOTA model robustness. We highlight the exceptional resilience of Claude-3.5-Sonnet in sharp contrast to the high vulnerability of Gemini-2.5-Pro. These findings show that scaling is a critical, measurable threat vector, and we urge the community to adopt large-scale robustness evaluation, rather than designing intricate algorithms on small ensembles, to build and validate genuinely robust and trustworthy foundation models.
@@ 2 Backgrounds
In this section, we provide necessary backgrounds, including the problem formulations for different attack settings and the prerequisite algorithms.
2.1 Problem Formulations
Our work investigates scaling laws across multiple black-box attack scenarios. We formulate the settings below.
2.1.1 Attacks on Image Classifiers
Let $\mathcal{F}$ denote the set of all image classifiers. The vulnerability of such classifiers to adversarial examples has been a long-standing area of research [szegedy2013intriguing, goodfellow2014explaining]. Given a natural image $\bm{x}_{\text{nat}}$ with its true label $\bm{y}_{\text{nat}}$, the goal of a transfer-based attack is to craft an adversarial example $\bm{x}$ within an $\ell_{\infty}$ perturbation budget $\epsilon$ that fools black-box target models.
Targeted Attack. The primary focus of our work is the challenging targeted attack setting. Given a target class $\bm{y}$ (where $\bm{y}\neq\bm{y}_{\text{nat}}$), the goal is to craft an $\bm{x}$ that is misclassified as $\bm{y}$. This can be formulated as minimizing the expected loss over all models in $\mathcal{F}$:
$\min_{\bm{x}}\mathbb{E}_{f\in\mathcal{F}}\left[\mathcal{L}(f(\bm{x}),\bm{y})\right],\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (1)
where $\mathcal{L}$ is a loss function (e.g., cross-entropy). In practice, we approximate this by ensembling a limited set of $T$ surrogate classifiers $\{f_{i}\}_{i=1}^{T}\subset\mathcal{F}$:
$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (2)
Untargeted Attack. We also validate our scaling laws in the untargeted setting in Sec. 5.1, where the goal is to mislead the classifier to any class other than $\bm{y}_{\text{nat}}$. The objective is to maximize the loss with respect to original $\bm{y}_{\text{nat}}$:
$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon.$ (3)
2.1.2 Attacks on Vision Encoders
Contrastive Language-Image Pretraining (CLIP) is a popular self-supervised framework that can pretrain large-scale language-vision models on web-scale text-image pairs via contrastive learning [radford2021learningtransferablevisualmodels, palatucci2009zero, lampert2009learning]. Based on the generalization ability of CLIP models, previous works have utilized them as surrogate models in model ensembles [dong2023robust, anonymous2024tuap, fort2024ensembleeverywheremultiscaleaggregation].
We extend our analysis to attack these vision encoders, which are foundational to modern MLLMs. Instead of classifying to a specific logit, the goal is to craft an image $\bm{x}$ whose embedding $f^{I}(\bm{x})$ is close to the text embedding $\bm{e}_{\text{tar}}$ of a target label $\bm{y}$. This target embedding is typically pre-computed using a text encoder, $f^{T}$. The objective function is adapted by replacing the classifier $f_{i}$ with a surrogate image encoder $f_{i}^{I}$ and the cross-entropy loss $\mathcal{L}$ with a similarity-based loss $\mathcal{L}_{\text{sim}}$ (e.g., cosine embedding loss):
$\min_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}_{\text{sim}}(f_{i}^{I}(\bm{x}),\bm{e}_{\text{tar}}),\quad\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (4)
where $\bm{e}_{\text{tar}}=f^{T}(\bm{y})$ is the fixed target text embedding.
@@ 2.2 Prerequisite Algorithms
Researchers have developed sophisticated algorithms to enhance adversarial transferability, which can be categorized into gradient optimizers, input transformations, and ensemble strategies.
2.2.1 Gradient Optimizers
These methods refine the gradient update step itself. A pioneering example is MI-FGSM [dong2018boosting], which stabilizes gradients by integrating a momentum term:
@@ 2.2.1 Gradient Optimizers
with update $\bm{x}_{t+1}=\bm{x}_{t}+\alpha\cdot\operatorname{sign}(\bm{g}_{t+1})$. Other methods like VMI-FGSM [wang2021enhancing] further refine this momentum concept.
2.2.2 Input Transformations
These methods apply data augmentation during the attack to find more generalizable adversarial patterns. Prominent examples include Diverse Inputs (DI) [xie2019improving] and Spectrum Simulation Attack (SSA) [long2022frequency]. SSA, a state-of-the-art method, applies spectrum transformations $\mathcal{T}$ and computes gradients on the transformed input $\mathcal{T}(\bm{x})$.
Figure 2: Comparisons between CWA and a naïve ensemble (MI-FGSM with logit averaging). As $T$ increases, the naïve method suffers optimization stagnation (gradient norm collapses) due to gradient conflict. Advanced methods like CWA, which align gradients, avoid this issue entirely.
2.2.3 Ensemble Strategies and Gradient Conflict
The above components can be applied to ensemble attacks. The most straightforward method, often used with MI-FGSM, is logit ensembling [dong2018boosting], where the optimizer computes gradients over the sum of all models’ logits.
While simple, this strategy exposes a critical flaw: gradient conflict. We discover that as the number of surrogates ($T$) increases, the gradients from different models, which often lack semantics and have random directions, tend to cancel each other out. As shown in Fig. 2, this conflict causes the ensemble gradient norm to collapse, leading to complete optimization stagnation.
To address this, advanced optimizers are required. These methods aim to find a common direction of vulnerability rather than simply averaging gradients. They typically work by reducing gradient variance, such as SVRE [xiong2022stochastic], or by encouraging gradient alignment, which is the approach taken by the Common Weakness Attack (CWA) [chen2023rethinking]. CWA seeks common vulnerabilities by encouraging flatter loss landscapes and minimizing the second-order Taylor series:
@@ 2.2.3 Ensemble Strategies and Gradient Conflict
where $\bm{p}_{i}$ is the closest optimum of model $f_{i}$ to $\bm{x}$ and $H_{i}$ is the Hessian. As Fig. 2 demonstrates, CWA effectively avoids the optimization stagnation seen in MI-FGSM.
Our Chosen Algorithm: SSA-CWA. For our main experiments, we require an algorithm that both resolves the gradient conflict and achieves state-of-the-art performance. We therefore select SSA-CWA, which combines the advanced input transformation of SSA [long2022frequency] with the robust gradient optimization of CWA [chen2023rethinking] (pseudocode provided in Appendix C). While leveraging SSA-CWA as our core algorithm, we also explore the scaling laws of other optimizers and validate our gradient conflict hypothesis in Sec. 5.1.
@@ 3 Scaling Laws for Adversarial Attacks
In this section, we present our key findings on scaling laws for black-box adversarial attacks. Sec. 3.1 introduces the theoretical motivation for our study. Sec. 3.2 details our experimental setup for generating adversarial examples. Sec. 3.3 presents the initial discovery of the empirical scaling laws on standard classifiers. Finally, Sec. 3.4 verifies the universality of these laws by testing them against state-of-the-art defended models and across diverse MLLMs.
3.1 Motivation
@@
Suppose that model ensemble $\{f_{i}\}_{i=1}^{T}$ is i.i.d. sampled from some distribution on $\mathcal{F}$, and $\hat{\bm{x}}$ goes to $\bm{x}^{*}$ as $T\to\infty$. Denote $\overset{d}{\to}$ as convergence in distribution, and $\mathrm{Cov}$ as the covariance matrix. Then we have the following asymptotic bound as $T\to\infty$:
$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}\mathrm{Cov}(\mathbb{L}(f_{i}))(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}),$ (9)
$T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\frac{1}{2}\|S\|_{2}^{2},\text{ with}$
@@
$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}\mathcal{N}(0,(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}),$ (10)
$T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\frac{1}{2}\|\mathcal{N}(0,\bm{I})\|_{2}^{2}.$
@@
Theorem 2 holds since $\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*})))=\nabla^{2}\mathbb{L}(\bm{x}^{*})$ for the cross-entropy loss. Therefore, we have $2T(\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*}))\overset{d}{\to}\chi^{2}(D)$ where $D$ is the dimension of the input, which means $\mathbb{L}(\hat{\bm{x}})-\mathbb{L}(\bm{x}^{*})$ converges in $\mathcal{O}(\frac{D}{2T})$.
These theorems provide the core theoretical motivation for our work: increasing the ensemble cardinality $T$ is a principled approach, as it provably closes the gap between the empirical loss and the true population loss. However, we must acknowledge a crucial gap between theory and practical setting of black-box attacks. The theorems rely on i.i.d. assumptions that are not met in practice but for a theoretical convenience. In practice, our surrogate models are typically pre-trained on similar datasets (e.g., ImageNet) and share strong architectural biases, making them highly correlated [ilyas2019adversarialexamplesbugsfeatures]. Consequently, the asymptotic $\mathcal{O}(1/T)$ convergence rate derived in Remark 2 is unlikely to hold.
Therefore, we posit that while the theory provides a strong qualitative motivation (i.e., scaling $T$ should improve transferability), it does not offer a quantitative prediction for the empirical scaling law. The central question of how ASR and loss actually behave as $T$ increases in a practical setting remains open. This motivates our large-scale empirical investigation in the following sections to discover the true scaling laws governing black-box attacks.
3.2 Experimental Setup